Kyutiedask10:

A training field is formed by joining a rectangle and two semicircles, as shown below. The rectangle is 99m long and 72m wide. Find the area of the training field. Use the value 3.14 for π, and do not round your answer. Be sure to include the correct unit in your answer.

Vocaloid:



Vocaloid:

if we take the two semi-circles off the side, we get a full circle with a radius of 36 (72/2) so we add area of rectangle + area of circle = answer area of rectangle = L*W area of circle = pi * r^2

Allison:

If your question is answered, or no longer needs an answer, or if it's incomplete, please close it. @Ultrilliam if you can close this.